OpenStudy (kimjafo):

Let f(x)= -5x-4 and g(x)=6x-7. Find f(x)+g(x)

OpenStudy (kimjafo):

A. -11x-11 B. x+11 C. -11x=3b D. x+3

563blackghost (563blackghost):

Since we know that \(\Large{f(x)=-5x-4}\) and \(\Large{g(x)=6x-7}\) we would be able to find \(\Large{f(x)=g(x)}\) by inputting these equations... \(\huge{(-5x-4)+(6x-7)}\) So combine like terms...

OpenStudy (kimjafo):

Thanks for the help, I keep forgetting how to do these

563blackghost (563blackghost):

No problem ^^

OpenStudy (kimjafo):

x-11?

563blackghost (563blackghost):

Yup thats what I got but what I dont understand is that that is not one of your answer choices or did you type B wrong??

OpenStudy (kimjafo):

Yea D: I typed it wrong! Sorry!

563blackghost (563blackghost):

Its alright ^^